solve a proportion for x. 2x-5/x+2=3x-5/3x-1
Answer by kiru_khandelwal(79)

2x-5/x+2 = 3x-5/3x-1
=> (2x-5)(3x-1) = (3x-5)(x+2)
=> 2x(3x-1) -5(3x-1) = 3x(x+2) -5(x+2)
=> 6x^2 -2x -15x + 5 = 3x^2 + 6x -5x -10
=> 6x^2 - 17x + 5 = 3x^2 + x - 10
=> 6x^2 - 3x^2 -17x -x +5 + 10 = 0
=> 3x^2 - 18x + 15 = 0
Taking 3 as common facotr
=> 3(x^2 - 6x + 5) = 0
=> x^2 - 6x + 5 = 0
=> x^2 -x -5x + 5 = 0
=> x(x-1) -5(x-1) =0
=> (x-5)(x-1)=0
=> According to the zero product rule
either x-5 =0 => x=5
or x-1=0 => x=1
so x = 5 or 1